# ForestSplats: Deformable transient field for Gaussian Splatting in the Wild

### v1

@@ Abstract
Recently, 3D Gaussian Splatting (3D-GS) has emerged, showing real-time rendering speeds and high-quality results in static scenes. Although 3D-GS shows effectiveness in static scenes, their performance significantly degrades in real-world environments due to transient objects, lighting variations, and diverse levels of occlusion. To tackle this, existing methods estimate occluders or transient elements by leveraging pre-trained models or integrating additional transient field pipelines. However, these methods still suffer from two defects: 1) Using semantic features from the Vision Foundation model (VFM) causes additional computational costs. 2) The transient field requires significant memory to handle transient elements with per-view Gaussians and struggles to define clear boundaries for occluders, solely relying on photometric errors. To address these problems, we propose ForestSplats, a novel approach that leverages the deformable transient field and a superpixel-aware mask to efficiently represent transient elements in the 2D scene across unconstrained image collections and effectively decompose static scenes from transient distractors without VFM. We designed the transient field to be deformable, capturing per-view transient elements. Furthermore, we introduce a superpixel-aware mask that clearly defines the boundaries of occluders by considering photometric errors and superpixels. Additionally, we propose uncertainty-aware densification to avoid generating Gaussians within the boundaries of occluders during densification. Through extensive experiments across several benchmark datasets, we demonstrate that ForestSplats outperforms existing methods without VFM and shows significant memory efficiency in representing transient elements.
@@ 1 Introduction
Novel view synthesis from a sparse set of captured 2D images is a challenging and fundamental problem in computer vision and graphics, which has recently gained popularity. Such advancements play a critical role in applications such as augmented/virtual reality (VR) [5, 4], autonomous driving [11, 45], and 3D content generation [13, 34]. Even though Neural Radiance Field (NeRF) [16] has shown impressive results and rendering efficiency, it suffers from slow rendering speeds. Due to this issue, 3D Gaussian Splatting (3D-GS) [6] recently has emerged as an explicit 3D representation that offers real-time rendering speeds. However, existing methods [33, 29] assume that 2D images are captured without transient elements, various levels of occluders, or appearance variations such as changing sky, weather, and illumination. Therefore, accurately and robustly reconstructing with unconstrained image collections remains a primary challenge due to many transient elements in real-world environments. Several works have proposed various solutions to tackle this challenge. Early NeRF-based methods [15, 1, 36] introduce per-image transient embedding or photometric errors to typically ignore transient elements from unconstrained images. NeRF-W [15] optimizes an uncertainty embedding to address distractors. CR-NeRF [38] proposes a cross-ray paradigm and grid sampling strategy for efficient appearance modeling. Although showing promising results, these methods [15, 1, 38] are computationally intensive and struggle with real-time rendering. In contrast, 3D-GS-based methods [35, 9, 22, 12, 31] usually utilize semantic features from Vision Foundation Model (VFM) to estimate precise distractors mask, showing fast rendering speed. Wild-GS [35] uses depth information from Depth Anything [37] and a hierarchical decomposition strategy to regularize transient elements and preserve geometry consistency. WildGaussians [9] also utilizes semantic features from DINOv2 [18] to handle occlusions. HybridGS [12] defines a transient field to explicitly decompose transient elements. Despite significant advancements, these methods still suffer from two limitations: Firstly, leveraging semantic features from the Vision Foundation Model (VFM) causes additional computational costs due to pre-computation. Secondly, the transient field uses significant memory to handle transient elements using Gaussians per image and struggles to clearly estimate boundaries of occluders due to solely relying on the photometric errors. Our motivation stems from the lack of appearance variation and multi-view consistency in transient elements. That is, we only need to consider transient elements in the 2D scene from each view of unconstrained image scenarios. To end, we propose ForestSplats, which leverages a deformable transient field and superpixel-aware mask to effectively decouple transient elements from static scenes, showing high-quality rendering. We design deformable transient field that efficiently represents transient elements in the 2D scene across unconstrained image collections, which are explicitly decomposed from static field and show memory efficiency. Furthermore, we introduce superpixel-aware masking, which effectively separates distractors from static scenes without a pre-trained VFM by considering the photometric errors and superpixels in the 2D scene. Additionally, we incorporate a multi-stage training scheme to enhance the superpixel-aware mask in accurately capturing transient elements, achieving distractor-free novel-view synthesis. Furthermore, we introduce uncertainty-aware densification for static field, thereby avoiding generating Gaussians within the boundaries of occluders, showing high-quality rendering. Through extensive experiments on several datasets, we demonstrate that ForestSplats effectively decomposes transient elements from static scenes, outperforming the state-of-the-art methods without VFM in novel view synthesis. Our primary contributions can be summarized as follows.
@@ 4.2 Superpixel-aware Mask
Existing methods [21, 42, 9, 22] primarily address outlier handling using Least Squares techniques or by extracting semantic features from the Vision Foundation Model (VFM). Inspired by these works, we propose a superpixel-aware mask and a multi-stage training scheme that effectively decompose transient distractors from static scenes. Our approach considers photometric errors and superpixels. Following CR-NeRF [38], we obtain the transient mask $\mathcal{M}_{O}$ from Unet $S_{\theta}$ through a learning-based approach, optimizing to minimize the training loss. However, directly using the transient mask $\mathcal{M}_{O}$ often struggles with imprecise boundaries of occluders, as shown in Fig. 3. To tackle this, we leverage the superpixel algorithm [28], which considers color and spatial information. Given an each view $I_{gt}$ partitioned into set of superpixels $\mathcal{S}=\{s_{1},...,s_{j}\}$, we compute the mask coverage ratio $\rho_{j}$ for each superpixel $s_{j}$ as follow:
@@
Method Low Occlusion Medium Occlusion High Occlusion Average
@@
RobustNeRF [21] 16.60 0.407 0.480 21.72 0.741 0.248 20.60 0.602 0.379 19.64 0.583 0.369
Gaussian Opacity Field [41] 20.54 0.662 0.178 19.39 0.719 0.231 17.81 0.578 0.430 19.24 0.656 0.280
3D-GS [6] 19.68 0.649 0.199 19.19 0.709 0.220 19.03 0.649 0.340 19.30 0.669 0.253
Mip-Splatting [40] 20.15 0.661 0.194 19.12 0.719 0.221 18.10 0.664 0.333 19.12 0.681 0.249
GS-W [42] 18.67 0.595 0.288 21.50 0.783 0.152 18.52 0.644 0.335 19.56 0.674 0.258
NeRF On-the-go [20] 20.63 0.661 0.191 22.31 0.780 0.130 22.19 0.753 0.169 21.71 0.731 0.163
WildGaussians [9] 20.62 0.658 0.235 22.80 0.811 0.092 23.03 0.771 0.172 22.15 0.756 0.167
DeSplat [31] 19.93 0.695 0.170 23.47 0.845 0.100 24.33 0.870 0.105 22.58 0.803 0.125
HybridGS [12] 21.42 0.684 0.268 23.51 0.830 0.160 23.05 0.768 0.204 22.66 0.761 0.211
SpotLessSplats [22] 22.67 0.785 0.165 24.33 0.880 0.100 24.30 0.865 0.140 23.76 0.843 0.135
ForestSplats (Ours) 21.89 0.738 0.216 23.96 0.870 0.135 24.34 0.815 0.119 23.40 0.807 0.156
@@ 5 Experiments
Implementation details. We develop our ForestSplats method based on WildGaussians [9]. We follow the hyperparameter settings for the default setting and train using the Adam [7] optimizer without weight decay. Specifically, we conducted the improved ADC solely for a static field. Furthermore, we extended the implementation by adapting the Pixel-GS [44] to calculate the number of per-pixel Gaussians using a custom CUDA kernel. For transient elements, we adopt the transient field initialized by randomly selecting 10K points from those obtained through SfM. The deformable MLP consists of eight fully connected layers. In the initial training phase, we conduct 150K steps of training solely on the static field to represent the entire image. Subsequently, we proceed to joint training, where both static and transient fields are optimized until 200K iterations. Finally, we focus on training the static field to capture the static scene. For a fair comparison, we use NerfBaselines [8] as our evaluation framework, providing a unified interface to the originally released source codes. All experiments are performed on an NVIDIA RTX 5000 GPU.
@@ 7 Ablation Study
As shown in Tab. 3, the deformable transient field significantly reduces the CPU memory usage (MB). Specifically in the Sacre Coeur, although image-per-time embedding is considered, the deformable transient field achieves over an 80% improvement in CPU memory efficiency (MB). Compared to existing methods [31, 12], the deformable transient field enables all transient Gaussians to be loaded onto the GPU, avoiding unnecessary context switching between the CPU and GPU. Furthermore, we also analyze the performance based on the number of transient Gaussians as shown in Fig. 7-(a). We observe that the number of transient Gaussians primarily aims to occupy the regions with transient elements, which results in a slight effect on the representation of the static field. Moreover, we compare the transient field with and without the deformation MLP, as shown in Tab. 4. These results indicate that the deformable transient field not only significantly reduces memory usage (MB) but also facilitates scalability, particularly in large-scale scenarios.
@@ 8 Limitation and Conclusion
Although the transient field does not exhibit high quality due to deformable Gaussians, as shown in Fig. 8, the deformable transient field effectively captures transient elements while showing significant memory efficiency. Another limitation of our forestsplats is its difficulty in generalizing the number of superpixels due to the varying levels of occlusion in datasets, as depicted in Fig. 9-(a).

### v2

@@ Abstract
Recently, 3D Gaussian Splatting (3D-GS) has emerged, showing real-time rendering speeds and high-quality results in static scenes. Although 3D-GS shows effectiveness in static scenes, their performance significantly degrades in real-world environments due to transient objects, lighting variations, and diverse levels of occlusion. To tackle this, existing methods estimate occluders or transient elements by leveraging pre-trained models or integrating additional transient field pipelines. However, these methods still suffer from two defects: 1) Using semantic features from the Vision Foundation model (VFM) causes additional computational costs due to pre-training on external data to use prior knowledge. 2) The transient field requires significant memory to handle transient elements with per-view Gaussians and struggles to define clear boundaries for occluders, solely relying on photometric errors. To address these problems, we propose ForestSplats, a novel approach that leverages the deformable transient field and a superpixel-aware mask to efficiently represent transient elements in the 2D scene across unconstrained image collections and effectively decompose static scenes from transient distractors without VFM. We designed the transient field to be deformable, capturing per-view transient elements. Furthermore, we introduce a superpixel-aware mask that clearly defines the boundaries of occluders by considering photometric errors and superpixels. Additionally, we propose uncertainty-aware densification to avoid generating Gaussians within the boundaries of occluders during densification. Through extensive experiments across several benchmark datasets, we demonstrate that ForestSplats outperforms existing methods without VFM and shows significant memory efficiency in representing transient elements.
@@ 1 Introduction
Novel view synthesis from a sparse set of captured 2D images is a challenging and fundamental problem in computer vision and graphics, which has recently gained popularity. Such advancements play a critical role in applications such as augmented/virtual reality (VR) [6, 5], autonomous driving [12, 47], and 3D content generation [15, 36]. Even though Neural Radiance Field (NeRF) [18] has shown impressive results and rendering efficiency, it suffers from slow rendering speeds. Due to this issue, 3D Gaussian Splatting (3D-GS) [7] recently has emerged as an explicit 3D representation that offers real-time rendering speeds. However, existing methods [35, 31] assume that 2D images are captured without transient elements, various levels of occluders, or appearance variations such as changing sky, weather, and illumination. Therefore, accurately and robustly reconstructing with unconstrained image collections remains a primary challenge due to many transient elements in real-world environments. Several works have proposed various solutions to tackle this challenge. Early NeRF-based methods [17, 2, 38] introduce per-image transient embedding or photometric errors to typically ignore transient elements from unconstrained images. NeRF-W [17] optimizes an uncertainty embedding to address distractors. CR-NeRF [40] proposes a cross-ray paradigm and grid sampling strategy for efficient appearance modeling. Although showing promising results, these methods [17, 2, 40] are computationally intensive and struggle with real-time rendering. In contrast, 3D-GS-based methods [37, 10, 24, 14, 33] usually utilize semantic features from Vision Foundation Model (VFM) to estimate precise distractors mask, showing fast rendering speed. Wild-GS [37] uses depth information from Depth Anything [39] and a hierarchical decomposition strategy to regularize transient elements and preserve geometry consistency. WildGaussians [10] also utilizes semantic features from DINOv2 [20] to handle occlusions. HybridGS [14] defines a transient field to explicitly decompose transient elements. Despite significant advancements, these methods still suffer from two limitations: Firstly, leveraging semantic features from the Vision Foundation Model (VFM) causes additional computational costs due to pre-computation on external huge data to use prior knowledge. Secondly, the transient field uses significant memory to handle transient elements using Gaussians per image and struggles to clearly estimate boundaries of occluders due to solely relying on the photometric errors. Our motivation stems from the lack of appearance variation and multi-view consistency in transient elements. That is, we only need to consider transient elements in the 2D scene from each view of unconstrained image scenarios. To end, we propose ForestSplats, which leverages a deformable transient field and superpixel-aware mask to effectively decouple transient elements from static scenes, showing high-quality rendering. We design deformable transient field that efficiently represents transient elements in the 2D scene across unconstrained image collections, which are explicitly decomposed from static field and show memory efficiency. Furthermore, we introduce superpixel-aware masking, which effectively separates distractors from static scenes without a pre-trained VFM by considering the photometric errors and superpixels in the 2D scene. Additionally, we incorporate a multi-stage training scheme to enhance the superpixel-aware mask in accurately capturing transient elements, achieving distractor-free novel-view synthesis. Furthermore, we introduce uncertainty-aware densification for static field, thereby avoiding generating Gaussians within the boundaries of occluders, showing high-quality rendering. Through extensive experiments on several datasets, we demonstrate that ForestSplats effectively decomposes transient elements from static scenes, outperforming the state-of-the-art methods without VFM in novel view synthesis. Our primary contributions can be summarized as follows.
@@ 4.2 Superpixel-aware Mask
Existing methods [23, 44, 10, 24] primarily address outlier handling using Least Squares techniques or by extracting semantic features from the Vision Foundation Model (VFM). Inspired by these works, we propose a superpixel-aware mask and a multi-stage training scheme that effectively decompose transient distractors from static scenes. Our approach considers photometric errors and superpixels. Following CR-NeRF [40], we obtain the transient mask $\mathcal{M}_{O}$ from Unet $S_{\theta}$ through a learning-based approach, optimizing to minimize the training loss without per-scene optimization. However, directly using the transient mask $\mathcal{M}_{O}$ often struggles with imprecise boundaries of occluders, as shown in Fig. 3. To tackle this, we leverage the superpixel algorithm [30], which considers color and spatial information. Given an each view $I_{gt}$ partitioned into set of superpixels $\mathcal{S}=\{s_{1},...,s_{j}\}$, we compute the mask coverage ratio $\rho_{j}$ for each superpixel $s_{j}$ as follow:
@@
Method GPU hrs / FPS Low Occlusion Medium Occlusion High Occlusion Average
@@
RobustNeRF [23] - / $<$1 16.60 0.407 0.480 21.72 0.741 0.248 20.60 0.602 0.379 19.64 0.583 0.369
Gaussian Opacity Field [43] 0.41 / 43 20.54 0.662 0.178 19.39 0.719 0.231 17.81 0.578 0.430 19.24 0.656 0.280
3D-GS [7] 0.35 / 116 19.68 0.649 0.199 19.19 0.709 0.220 19.03 0.649 0.340 19.30 0.669 0.253
Mip-Splatting [42] 0.18 / 82 20.15 0.661 0.194 19.12 0.719 0.221 18.10 0.664 0.333 19.12 0.681 0.249
GS-W [44] 0.55 / 71 18.67 0.595 0.288 21.50 0.783 0.152 18.52 0.644 0.335 19.56 0.674 0.258
NeRF On-the-go [22] 43 / $<$1 20.63 0.661 0.191 22.31 0.780 0.130 22.19 0.753 0.169 21.71 0.731 0.163
WildGaussians [10] 0.50 / 108 20.62 0.658 0.235 22.80 0.811 0.092 23.03 0.771 0.172 22.15 0.756 0.167
SpotLessSplats [24] 0.34 / 93 20.01 0.596 0.276 22.79 0.816 0.161 21.91 0.71 0.222 21.57 0.707 0.219
DeSplat [33] 0.24 / 106 19.93 0.695 0.170 23.47 0.845 0.100 24.33 0.870 0.105 22.58 0.803 0.125
HybridGS [14] - 21.42 0.684 0.268 23.51 0.830 0.160 23.05 0.768 0.204 22.66 0.761 0.211
ForestSplats (Ours) 1.13 / 117 21.89 0.738 0.216 23.96 0.870 0.135 24.34 0.815 0.119 23.40 0.807 0.156
@@ 5 Experiments
11footnotetext: The results are reproduced from the official code.
@@
Implementation details. We develop our ForestSplats method based on WildGaussians [10]. We follow the hyperparameter settings for the default setting and train using the Adam [8] optimizer without weight decay. Specifically, we conducted the improved ADC solely for a static field. Furthermore, we extended the implementation by adapting the Pixel-GS [46] to calculate the number of per-pixel Gaussians using a custom CUDA kernel. For transient elements, we adopt the transient field initialized by randomly selecting 10K points from those obtained through SfM. The deformable MLP consists of eight fully connected layers. In the initial training phase, we conduct 150K steps of training solely on the static field to represent the entire image. Subsequently, we proceed to joint training, where both static and transient fields are optimized until 200K iterations. Finally, we focus on training the static field to capture the static scene. For a fair comparison, we use NerfBaselines [9] as our evaluation framework, providing a unified interface to the originally released source codes. All experiments are performed on an NVIDIA RTX 24GB GPU.
@@ 7 Ablation Study
As shown in Tab. 3, the deformable transient field significantly reduces the CPU memory usage (MB). Specifically in the Sacre Coeur, although transient embedding per image is considered, the deformable transient field achieves over an 80% improvement in CPU memory efficiency (MB). Compared to existing methods [33, 14], the deformable transient field enables all transient Gaussians to be loaded onto the GPU, avoiding unnecessary context switching between the CPU and GPU. Furthermore, we also analyze the performance based on the number of transient Gaussians as shown in Fig. 7-(a). We observe that the number of transient Gaussians primarily aims to occupy the regions with transient elements, which results in a slight effect on the representation of the static field. Moreover, we compare the transient field with and without the deformation MLP, as shown in Tab. 4. These results indicate that the deformable transient field not only significantly reduces memory usage (MB) but also facilitates scalability, particularly in large-scale scenarios.
@@ 8 Limitation and Conclusion
Although the transient field does not exhibit high quality due to deformable Gaussians, as shown in Fig. 8, the deformable transient field effectively captures transient elements while showing significant memory efficiency. Another limitation of our forestsplats is its difficulty in generalizing the number of superpixels due to the varying levels of occlusion in datasets, as depicted in Fig. 9.